# JAX-MPM: A Learning-Augmented Differentiable Meshfree Framework for GPU-Accelerated Lagrangian Simulation and Geophysical Inverse Modeling

### v1

@@ Abstract
Differentiable programming that enables automatic differentiation through simulation pipelines has emerged a powerful paradigm in scientific computing, naturally supporting forward and inverse modeling and facilitating seamless integration with deep learning frameworks. In this study, we present JAX-MPM, a general-purpose differentiable meshfree solver within a hybrid Lagrangian–Eulerian framework, tailored specifically for simulating complex continuum mechanics problems involving large deformations, frictional contact, and inelastic material behavior, with particular emphasis on geomechanics and geophysical hazard applications. The proposed theoretical framework is built upon the material point method (MPM) and implemented in the modern JAX computing architecture, resulting in a fully differentiable simulator. Leveraging high-performance GPU computing, this framework supports efficient gradient-based optimization directly through its time-stepping solvers, enabling high-fidelity forward and inverse modeling of large-scale continuum systems. Moreover, JAX-MPM inherently supports the joint training of physical models and neural networks, allowing the learning of embedded hidden system closures or constitutive models parameterized by neural networks. We validate the JAX-MPM framework through a series of 2D and 3D benchmark simulations, including dam-breaks and granular collapses, demonstrating its numerical accuracy and GPU-accelerated performance. We show that a high-resolution 3D simulation of a granular cylinder collapse with 2.7 million particles completes 1000 time steps in around 22 seconds (single precision) and 98 seconds (double precision) on a single GPU. Beyond forward modeling, we demonstrate inverse modeling capabilities through applications such as velocity field reconstruction and spatially varying friction estimation. Collectively, these results establish JAX-MPM as a unified and scalable platform for differentiable meshfree modeling, advancing data assimilation and physical inference for complex solid and geophysical applications.
@@
Differentiable programming , Lagrangian meshfree formulation , MPM , Geophysical hazards , Inverse modeling , Scientific machine learning , JAX
journal:
@@ 1.1 Overview
Natural hazards or geohazards [36, 46]—such as landslides, dam-breaks, debris flows, and soil liquefaction —are commonly recognized as gravitationally-driven processes, which involve large-scale failure of geomaterials (e.g., soils, rocks, and granular media that constitute the Earth’s near-surface), posing substantial risks to both natural and built environments. These phenomena are governed by highly nonlinear, path-dependent material behavior [95], and typically involve severe deformations and rapidly evolving boundary conditions [4], making predictive high-fidelity modeling and simulation of such hazard processes extremely challenging.
Numerical simulation has become a cornerstone of computational geomechanics, playing a critical role in geotechnical prediction and geohazard mitigation by enabling the analysis of complex geophysical processes primarily governed by partial differential equations (PDEs). Mesh-based methods such as the finite element method (FEM) [76, 61, 93] and finite volume method (FVM) [30, 29] have been widely adopted for this purpose. For example, FEM has been applied to simulate flow-like landslides [24], while FVM has been used to model coupled flow and shear failure in fractured geomechanical systems [29]. These methods have provided valuable insights in geohazard research and remain effective for problems involving moderate deformation and well-defined boundaries. However, in geohazard modeling scenarios characterized by extreme deformation, fragmentation, and evolving topologies, mesh-based methods often face numerical challenges due to severe mesh distortion and tangling, and can become unreliable or computationally impractical. To address these limitations, meshfree methods have been developed, representing materials as collections of particles or discrete points—eliminating the need for conforming meshes and improving robustness in large-deformation problems.
@@ 1.2 Lagrangian meshfree simulation
Among these methods, the Material Point Method (MPM) stands out due to its hybrid Eulerian–Lagrangian formulation. In this approach, material points carry internal state variables and track continuum deformation in a Lagrangian manner, while computational operations are performed on a background Eulerian grid that is reset at each time step. This decoupling of material and computational domains offers notable advantages [28], including avoiding mesh distortion and efficiently handling complex contact and costly particle–particle interactions via the background grid. These attributes make MPM particularly attractive for modeling challenging geomechanical problems [8, 3, 5] such as landslides, debris flows, and multiphase failure, where large deformations, fragmentation, and multi-body interactions are dominant. Given these strengths, MPM is adopted as the computational foundation for the differentiable simulation framework developed in this study.
@@ 1.3.2 Discrete approach: differentiable physics solver
In this work, we focus specifically on the differentiable implementation of MPM, which is inherently well-suited for modeling large deformation processes in geophysical systems. Traditional MPM implementations are typically developed in imperative programming languages such as C++ [60] , Fortran [102] , or high-level scientific computing platforms like Julia [86, 53]. While these implementations are efficient and widely used for forward simulations, they are not inherently designed to support differentiability or seamless integration with AD-based deep learning frameworks. Recent efforts have explored differentiable MPM solvers implemented in Taichi [50], a domain-specific programming language featuring built-in AD capacities [49]. Taichi has been successfully employed in various applications, including physics-based modeling and computer graphics [50, 49, 85], as well in tasks involving parameter and system identification [64, 97]. Despite these advances in MPM implementations, their development specifically tailored for large-scale, differentiable geomechanical modeling remains rare. Key limitations of existing approaches include restricted autodiff flexibility, limited multi-GPU support, and a separation between simulation and machine learning components.
@@ 1.4 A learning-based meshfree computational framework
To address this gap, we propose a general-purpose, Lagrangian meshfree solver built on JAX [14], a modern computational backend specifically tailored for differentiable programming. JAX integrates automatic differentiation, just-in-time (JIT) compilation, and GPU/TPU acceleration through the Accelerated Linear Algebra (XLA) compiler. Unlike imperative programming frameworks, which typically require explicit control over memory and execution flow, JAX promotes a functional, vectorized programming style, enabling efficient, loop-free simulations fully compatible with gradient-based optimization. Additionally, JAX integrates seamlessly with modern deep learning libraries such as Flax [43], Haiku [44], and Equinox [56], supporting the joint training of physical models and neural networks. JAX’s built-in parallelization tools, such as vmap and pmap, further facilitate efficient multi-GPU computation and batched operations, essential for large-scale inverse modeling and data-driven geomechanical workflows. Recently, JAX has been effectively combined with canonical numerical methods, such as FEM [98], finite difference method [69], finite-volume CFD solver [11], SPH [92] and meshless local Petrov-Galerkin method [32], among others.
Leveraging these capabilities, we develop JAX-MPM in this study, a GPU-accelerated, fully differentiable Material Point Method framework for large-deformation simulation and inverse modeling in geomechanics and solid mechanics. Implemented entirely in JAX, it unifies simulation, automatic differentiation, and neural networks within a single system, eliminating the need for external toolchains or inter-language bindings required by Taichi-based approaches. By combining the physical rigor of MPM with JAX’s flexibility, JAX-MPM enables robust, end-to-end pipelines for forward simulation, gradient-based inversion, and learning-driven optimization in natural hazard analysis and related applications.
@@
For inverse modeling, JAX-MPM introduces a unified PDE-constrained optimization framework that accommodates both Lagrangian (particle-based) and Eulerian (region-based) supervision without modifying the simulation pipeline. This flexible formulation allows integration of diverse observational inputs, such as sparse particle trajectories or fixed-location velocity sensors. We demonstrate its effectiveness on representative inverse tasks including initial velocity reconstruction and spatially varying friction estimation. These tasks span up to $\mathcal{O}(10^{3})$ trainable parameters and $\mathcal{O}(10^{4})$ degrees of freedom. In all cases, JAX-MPM achieves accurate, physically consistent reconstructions using only partial observational data, underscoring its potential for real-world geohazard applications.
@@ 2.1 Governing equations
$&\rho\dot{\bm{v}}=\nabla\cdot\bm{\sigma}+\rho\bm{b},\\ &\bm{\sigma}\cdot\bm{n}=\bar{\bm{t}}\quad\text{on }\Gamma_{N},\\ &\bm{u}=\bar{\bm{u}}\quad\text{on }\Gamma_{D},\end{aligned}\right.$ (1)
where $\rho$ denotes the material density, $\bm{b}$ is the body force per unit mass, and $\bm{\sigma}$ is the Cauchy stress tensor. The boundary $\partial\Omega$ is partitioned into Dirichlet ($\Gamma_{D}$) and Neumann ($\Gamma_{N}$) portions, corresponding to prescribed displacements and tractions, respectively. The velocity field is defined as $\bm{v}=\dot{\bm{u}}$, where $\bm{u}$ is the displacement field. Material time derivatives relate the kinematic quantities by $\bm{a}=\dot{\bm{v}}=\ddot{\bm{u}}$, with all derivatives taken following the material motion.
The weak form of Eq. (1) is derived by multiplying with a test function $\bm{w}$, integrating over the current configuration, and applying the divergence theorem:
$\int_{\Omega}\bm{w}\cdot\rho\bm{a}\mathrm{~{}d}\Omega=-\int_{\Omega}\nabla\bm{w}:\bm{\sigma}\mathrm{~{}d}\Omega+\int_{\Omega}\bm{w}\cdot\rho\bm{b}\mathrm{~{}d}\Omega+\int_{\Gamma_{{N}}}\bm{w}\cdot\bar{\bm{t}}\mathrm{~{}d}\Gamma$ (2)
@@
Following the standard Material Point Method (MPM) framework [90, 91], the continuum domain is discretized using a set of material points (also referred to as particles) located at positions $\bm{x}_{p}$, for $p=1,\dots,N_{p}$, which move freely over a fixed background Cartesian grid (illustrated in Fig. 1). Each material point represents a subdomain $\Omega_{p}\subset\Omega$, and carries the physical properties of the continuum, such as mass, volume, velocity, and stress. This particle-based discretization allows us to rewrite the weak form of the momentum balance equation (2) as a sum over all material points:
$\sum_{p}\int_{\Omega_{p}}\bm{w}\cdot\rho\bm{a}\mathrm{~{}d}\Omega=-\sum_{p}\int_{\Omega_{p}}\nabla\bm{w}:\bm{\sigma}\mathrm{~{}d}\Omega+\sum_{p}\int_{\Omega_{p}}\bm{w}\cdot\rho\bm{b}\mathrm{~{}d}\Omega+\sum_{p}\int_{\Gamma_{{N}}}\bm{w}\cdot\overline{\bm{t}}\mathrm{~{}d}\Gamma$ (3)
where the index $p$ denotes the material points in the domain. Note that the conservation of mass is inherently satisfied in the MPM formulation and is therefore omitted from this derivation [90].
To interpolate field variables onto the background grid, we introduce finite element (FE)-based nodal interpolation:
$\bm{a}=\sum_{i}\phi_{i}(\bm{x})\bm{a}_{i},\quad\bm{w}=\sum_{i}\phi_{i}(\bm{x})\bm{w}_{i}$ (4)
where the index $i$ denotes summation over over all $N_{g}$ grid nodes, and $\phi_{i}(\bm{x})$ is the shape function associated with node $i$. In this study, we employ a quadratic B-spline basis function $w(\xi)$ [88], defined as
$\omega(\xi)$ =$=\begin{cases}-|\xi|^{2}+\frac{3}{4}&0\leq|\xi|<\frac{1}{2}\\ \frac{1}{2}|\xi|^{2}-\frac{3}{2}|\xi|+\frac{9}{8}&\frac{1}{2}\leq|\xi|<\frac{3}{2}\\ 0&\frac{3}{2}\leq|\xi|\end{cases}$ (5)
The multidimensional shape function $\phi_{i}(\bm{x})$ is constructed using a tensor product of one-dimensional B-splines:
$\phi_{i}(\bm{x})=\prod_{\alpha=1}^{d}\omega\left(\frac{x_{\alpha}-x_{i\alpha}}{\Delta h}\right)$ (6)
where $\Delta h$ is the uniform grid spacing.
Substituting Eq. (4) into the discrete weak form and considering the arbitrariness of $\bm{w}_{i}$, we obtain the semi-discrete form:
$\sum_{j=1}^{N_{g}}m_{ij}\bm{a}_{j}=\bm{f}_{i}^{\mathrm{int}}+\bm{f}_{i}^{\mathrm{ext}}$ (7)
@@
$\bm{f}_{i}^{\mathrm{ext}}$ $=\sum_{p}\left[\int_{\Omega_{p}}\phi_{i}\rho\bm{b}\mathrm{~{}d}\Omega+\int_{\Gamma_{N}}\phi_{i}\bar{\bm{t}}\mathrm{~{}d}\Gamma\right]=\sum_{p}\phi_{ip}m_{p}\bm{b}_{p}+\bar{\bm{f}}_{i}$
where $\phi_{ip}=\phi_{i}(\bm{x}_{p})$, and $m_{p}$ is the mass of particle $p$. $m_{ij}$ is the consistent mass matrix. In this study, we adopt a lumped mass matrix, i.e., $m_{i}\equiv m_{ii}=\sum_{p}m_{p}\phi_{ip}$, to avoid an expensive matrix inversion [90]. Applying explicit time discretization, the semi-discrete momentum equation (8) becomes:
@@
where forward Euler integration is used for simplicity and efficiency. Here, $\bm{f}_{i}^{n}=\bm{f}_{i}^{\mathrm{int},n}+\bm{f}_{i}^{\mathrm{ext},n}$ is the total force acting on node $i$ at time step $n$, consisting of internal and external contributions. The superscript $n$ indicates the current time step, and $\Delta t=t^{n+1}-t^{n}$ is the time increment. To ensure the numerical stability of the explicit scheme, $\Delta t$ is determined according to the Courant–Friedrichs–Lewy (CFL) condition [23].

### v2

@@ Abstract
Differentiable programming has emerged as a powerful paradigm in scientific computing, enabling automatic differentiation through simulation pipelines and naturally supporting both forward and inverse modeling. We present JAX-MPM, a general-purpose differentiable meshfree solver based on the material point method (MPM) and implemented in the modern JAX architecture. The solver adopts a hybrid Eulerian–Lagrangian framework to capture large deformations, frictional contact, and inelastic material behavior, with emphasis on geomechanics and geophysical hazard applications. Leveraging GPU acceleration and automatic differentiation, JAX-MPM enables efficient gradient-based optimization directly through its time-stepping solvers and supports joint training of physical models with deep learning to infer unknown system conditions and uncover hidden constitutive parameters. We validate JAX-MPM through a series of 2D and 3D benchmark simulations, including dam-break and granular collapse problems, demonstrating both numerical accuracy and GPU-accelerated performance. Results show that a high-resolution 3D granular cylinder collapse with 2.7 million particles completes 1000 time steps in approximately 22 seconds (single precision) and 98 seconds (double precision) on a single GPU. Beyond high-fidelity forward modeling, we demonstrate the framework’s inverse modeling capabilities through tasks such as velocity field reconstruction and the estimation of spatially varying friction from sparse data. In particular, JAX-MPM accommodates data assimilation from both Lagrangian (particle-based) and Eulerian (region-based) observations, and can be seamlessly coupled with neural network representations. These results establish JAX-MPM as a unified and scalable differentiable meshfree platform, that advances fast physical simulation and data assimilation for complex solid and geophysical systems.
@@
Differentiable simulator , Meshfree Lagrangian formulation , MPM , Geophysical hazards , Inverse modeling , Scientific machine learning , JAX
@@ 1.1 Overview
Natural hazards, or geohazards [36, 46], such as landslides, dam-break flows, debris flows, and soil liquefaction, are gravitationally driven processes involving large-scale failure of geomaterials (e.g., soils, rocks, and granular media in the Earth’s near-surface) and they pose significant risks to both natural and built environments. Governed by highly nonlinear, path-dependent material behavior [94], these phenomena often entail severe deformations and rapidly evolving boundary conditions [4], making high-fidelity predictive modeling and simulation particularly challenging.
Numerical modeling and simulation are essential tools in geomechanics and geotechnical engineering, enabling the analysis of complex physical processes through the solution of partial differential equations (PDEs). Mesh-based methods such as the finite element method (FEM) [75, 60, 92] and finite volume method (FVM) [29, 28] have been widely adopted for this purpose. For example, FEM has been applied to simulate flow-like landslides [23], while FVM has been used to model coupled flow and shear failure in fractured geomechanical systems [28]. These methods have provided valuable insights in geohazard research and remain effective for problems involving moderate deformation and well-defined boundaries. However, in geohazard modeling scenarios characterized by extreme deformation, fragmentation, and evolving topologies, mesh-based methods often face numerical challenges due to severe mesh distortion and tangling, and can become unreliable or computationally impractical. To address these limitations, meshfree methods have been developed, representing materials as collections of particles or discrete points—eliminating the need for conforming meshes and improving robustness in large-deformation problems.
@@ 1.2 Lagrangian meshfree simulation
Among these methods, MPM stands out due to its hybrid Eulerian–Lagrangian formulation. In this approach, material points carry internal state variables and track continuum deformation in a Lagrangian manner, while computational operations are performed on a background Eulerian grid that is reset at each time step. This decoupling of material and computational domains offers notable advantages [27], including avoiding mesh distortion and efficiently handling complex contact and costly particle–particle interactions via the background grid. These attributes make MPM particularly attractive for modeling challenging geomechanical problems [8, 3, 5] such as landslides, debris flows, and multiphase failure, where large deformations, fragmentation, and multi-body interactions are dominant. Given these strengths, MPM is adopted as the computational foundation for the differentiable simulation framework developed in this study.
@@ 1.3.2 Discrete approach: differentiable physics solver
In this work, we focus specifically on the differentiable implementation of MPM, which is inherently well-suited for modeling large deformation processes in geophysical systems. Traditional MPM implementations are typically developed in imperative programming languages such as C++ [59], Fortran [101], or high-level scientific computing platforms like Julia [85, 52]. While these implementations are efficient and widely used for forward simulations, they are not inherently designed to support differentiability or seamless integration with AD-based deep learning frameworks. Recent efforts have explored differentiable MPM solvers implemented in Taichi [49], a domain-specific programming language featuring built-in AD capacities [48]. Taichi has been successfully employed in various applications, including physics-based modeling and computer graphics [49, 48, 84], as well as in tasks involving parameter and system identification [63, 96]. Despite these advances in MPM implementations, their development specifically tailored for large-scale, differentiable geomechanical modeling remains rare. Key limitations of existing approaches include restricted AD flexibility, limited multi-GPU support, and a separation between simulation and machine learning components.
@@ 1.4 A learning-based meshfree computational framework
To address this gap, we propose a general-purpose, Lagrangian meshfree solver built on JAX [13], a modern computational backend specifically tailored for differentiable programming. JAX integrates automatic differentiation, just-in-time (JIT) compilation, and GPU/TPU acceleration through the accelerated linear algebra (XLA) compiler. Unlike imperative programming frameworks, which typically require explicit control over memory and execution flow, JAX promotes a functional, vectorized programming style, enabling efficient, loop-free simulations fully compatible with gradient-based optimization. Additionally, JAX integrates seamlessly with modern deep learning libraries such as Flax [43], Haiku [44], and Equinox [55], supporting the joint training of physical models and neural networks. JAX’s built-in parallelization tools, such as vmap and pmap, further facilitate efficient multi-GPU computation and batched operations, which are essential for large-scale inverse modeling and data-driven geomechanical workflows. Recently, JAX has been effectively combined with canonical numerical methods, such as FEM [97], finite difference method [68], finite-volume CFD solver [10], SPH [91] and meshless local Petrov-Galerkin method [31], among others.
Leveraging these capabilities, we develop JAX-MPM in this study, a GPU-accelerated, fully differentiable meshfree framework for large-deformation simulation and inverse modeling in geomechanics and solid mechanics. Implemented entirely in JAX, it unifies simulation, automatic differentiation, and neural networks within a single system, eliminating the need for external toolchains or inter-language bindings required by Taichi-based approaches. By combining the physical rigor of MPM with JAX’s flexibility, JAX-MPM enables robust, end-to-end pipelines for forward simulation, gradient-based inversion, and learning-driven optimization in natural hazard analysis and related applications.
@@
For inverse modeling, JAX-MPM introduces a unified PDE-constrained optimization framework that accommodates both Lagrangian (particle-based) and Eulerian (region-based) supervision without modifying the simulation pipeline. This flexible formulation allows the integration of diverse observational inputs, such as sparse particle trajectories or fixed-location velocity sensors. We demonstrate its effectiveness on representative inverse tasks including initial velocity reconstruction and spatially varying friction estimation. These tasks span up to $\mathcal{O}(10^{3})$ trainable parameters and $\mathcal{O}(10^{4})$ degrees of freedom. In all cases, JAX-MPM achieves accurate, physically consistent reconstructions using only partial observational data, underscoring its potential for real-world geohazard applications. To the best of our knowledge, this is the first attempt to employ a differentiable Lagrangian physics solver for large-scale inverse modeling in geophysical flow applications.
@@ 2.1 Governing equations
$\left\{\begin{aligned} &\rho\dot{\bm{v}}=\nabla\cdot\bm{\sigma}+\rho\bm{b},\\ &\bm{\sigma}\cdot\bm{n}=\bar{\bm{t}}\quad\text{on }\Gamma_{N},\\ &\bm{u}=\bar{\bm{u}}\quad\text{on }\Gamma_{D},\end{aligned}\right$. (1)
where $\rho$ denotes the material density, $\bm{b}$ is the body force per unit mass, and $\bm{\sigma}$ is the Cauchy stress tensor. The boundary $\partial\Omega$ is partitioned into Dirichlet ($\Gamma_{D}$) and Neumann ($\Gamma_{N}$) portions, corresponding to prescribed displacements $\bar{\bm{u}}$ and tractions $\bar{\bm{t}}$, respectively. The velocity field is defined as $\bm{v}=\dot{\bm{u}}$, where $\bm{u}$ is the displacement field, and the superposed dot denotes the material derivative.
The weak form of Eq. (1) is derived by multiplying with a vector-valued test function $\bm{w}$, integrating over the current configuration, and applying the divergence theorem:
$\int_{\Omega}\bm{w}\cdot\rho\dot{\bm{v}}\mathrm{~d}\Omega=-\int_{\Omega}\nabla\bm{w}:\bm{\sigma}\mathrm{~d}\Omega+\int_{\Omega}\bm{w}\cdot\rho\bm{b}\mathrm{~d}\Omega+\int_{\Gamma_{{N}}}\bm{w}\cdot\bar{\bm{t}}\mathrm{~d}\Gamma$ (2)
@@
Following the standard MPM framework [89, 90] (illustrated in Fig. 1), the continuum domain is discretized into a collection of Lagrangian material points (particles) located at positions $\bm{x}_{p}$, $p=1,\dots,N_{p}$. These particles move freely over a fixed Eulerian background grid, whose nodes are located at $\bm{x}_{i}$, $i=1,\dots,N_{g}$. Throughout this paper, the subscript $p$ refers to material points, while the subscript $i$ refers to grid nodes.
Each material point is associated with a subdomain $\Omega_{p}\subset\Omega$, and it carries the physical properties of the continuum, such as mass, volume, velocity, and stress. This particle-based representation allows the weak form of the momentum balance equation (2) to be expressed as a sum over all material points:
$\sum_{p}\int_{\Omega_{p}}\bm{w}\cdot\rho\dot{\bm{v}}\mathrm{~d}\Omega=-\sum_{p}\int_{\Omega_{p}}\nabla\bm{w}:\bm{\sigma}\mathrm{~d}\Omega+\sum_{p}\int_{\Omega_{p}}\bm{w}\cdot\rho\bm{b}\mathrm{~d}\Omega+\sum_{p}\int_{\Gamma_{{N},p}}\bm{w}\cdot\overline{\bm{t}}\mathrm{~d}\Gamma$ (3)
where the index $p$ denotes the material points in the domain, and $\Gamma_{{N},p}$ denotes the portion of the Neumann boundary $\Gamma_{N}$ intersecting the subdomain $\Omega_{p}$. It is worth noting that the conservation of mass is inherently satisfied in the MPM formulation and is therefore omitted from the present derivation [89].
The velocity $\bm{v}(\bm{x})$ and test function $\bm{w}(\bm{x})$ are discretized on the background grid as linear combinations of the nodal shape functions $\phi_{i}(\bm{x})$:
$\bm{v}=\sum_{i}\phi_{i}(\bm{x})\bm{v}_{i},\quad\bm{w}=\sum_{i}\phi_{i}(\bm{x})\bm{w}_{i}$ (4)
where the index $i=1,...,N_{g}$ denotes the grid nodes, $\phi_{i}(\bm{x})$ denotes the shape function associated with node $i$ with a compact support, and $\bm{v}_{i}$ and $\bm{w}_{i}$ are the corresponding nodal velocity and nodal test function, respectively. In this study, the multidimensional shape function $\phi_{i}(\bm{x})$ is constructed using a tensor product of one-dimensional B-splines:
$\phi_{i}(\bm{x})=\prod_{\alpha=1}^{d}\bar{\phi}\left(\frac{x_{\alpha}-x_{i\alpha}}{\Delta h}\right)$ (5)
where $\Delta h$ is the uniform grid spacing, and the quadratic B-spline basis function $\bar{\phi}$ is defined as [87]:
[…]
Substituting Eq. (4) into the discrete weak form (3) and considering the arbitrariness of nodal test functions $\bm{w}_{i}$, we obtain the semi-discrete form:
$\sum_{j=1}^{N_{g}}m_{ij}\dot{\bm{v}}_{j}=\bm{f}_{i}^{\mathrm{int}}+\bm{f}_{i}^{\mathrm{ext}}$ (7)
@@
$\bm{f}_{i}^{\mathrm{ext}}$ $=\sum_{p}\left[\int_{\Omega_{p}}\phi_{i}\rho\bm{b}\mathrm{~d}\Omega\right]+\sum_{p}\int_{\Gamma_{{N},p}}\phi_{i}\bar{\bm{t}}\mathrm{~d}\Gamma=\sum_{p}\phi_{ip}m_{p}\bm{b}_{p}+\bar{\bm{f}}_{i}^{t}$
where $\phi_{ip}=\phi_{i}(\bm{x}_{p})$, $V_{p}$, $m_{p}$ and $\bm{\sigma}_{p}$ are the volume, mass and stress tensor associated with particle $p$, respectively, $m_{ij}$ is the consistent mass matrix, and $\bar{\bm{f}}_{i}^{t}$ is the nodal force contributed by the traction $\bar{\bm{t}}$.
To avoid the costly inversion of the consistent mass matrix [89], we adopt a lumped mass matrix, which retains only the nonzero diagonal entries, $m_{i}=\sum_{p}m_{p}\phi_{ip}$. With this lumped mass matrix, application of forward Euler time integration to the semi-discrete momentum equation (7) yields the following explicit time-stepping scheme:
@@
where the superscript $n$ denotes the current time step, $\Delta t=t^{n+1}-t^{n}$ is the time increment, and $\bm{f}_{i}^{n}=\bm{f}_{i}^{\mathrm{int},n}+\bm{f}_{i}^{\mathrm{ext},n}$ is the total force acting on node $i$, consisting of internal and external contributions. To ensure the numerical stability of the explicit scheme, $\Delta t$ is chosen according to the Courant–Friedrichs–Lewy (CFL) condition [22].
Figure 1: Illustration of the material point method (MPM) procedure. The procedure consists of three stages: P2G transfer, grid-based updates, G2P transfer. The G2P operation includes grid-to-particle interpolation, particle-based updates, and grid reset. Black dots denote the fixed background grid, blue stars represent particles, and purple stars indicate updated particles. Dashed orange arrows illustrate interpolation contributions, while solid black arrows indicate velocity vectors.